OpenStudy (anonymous):

1. Simplify: (2x + 9)(4x2 + 2x - 12)

OpenStudy (anonymous):

Distribution :) Just do it :> a(p+q+r) = ap + aq + ar

OpenStudy (anonymous):

oh! and that 4x2 is actually supposed to be a 4x^2

OpenStudy (anonymous):

But how would i be able to use distribution on this problem when there is not a number on the beginning of the outside of theses expressions?

OpenStudy (anonymous):

Don't worry, I know :) I got ya :3 so... distribute (2x + 9)(4x² + 2x - 12) = 4x²(2x+9) + 2x(2x+9) - 12(2x+9) It may not be A number, but it can still be distributed, as a whole :)

OpenStudy (anonymous):

Oh okay! So I would solve the same way I would solve a regular distribution problem?

OpenStudy (anonymous):

Yeah :) But now, you have to distribute again :)

OpenStudy (anonymous):

hold on I have to do this on paper! lol I am really slow at math!

OpenStudy (anonymous):

Take your time :3

OpenStudy (anonymous):

@PeterPan would you also multiply 4x^2 by the +9?

OpenStudy (anonymous):

Of course :) It's a distribution just like any other ^.^

OpenStudy (anonymous):

Oh okay! So @PeterPan when you multiply the 4x twice you get 16 or would you just leave that 4x^2?

OpenStudy (anonymous):

Let's do that first part :) 4x²(2x+9) We multiply 4x² to 2x, we get...8x³ We multiply 4x² to 9, we get 36x² So... 4x²(2x+9) = 8x³ + 36x² Any questions? Do the rest ^.^

OpenStudy (anonymous):

well isnt 4x times 4x 16? because dont you have to multiply it by the same number twice not by 2? because thats how my teacher explained it.

OpenStudy (anonymous):

Of course 4 x 4 = 16 But why would this interest us? There is no instance of 4x4 in the problem :)

OpenStudy (anonymous):

yeah, but i meant that since their 4x to the power of 2 wouldnt that mean it would be 4x multiplied by 4x?

OpenStudy (anonymous):

AHH \[\huge 4x^2=4\cdot x\cdot x \ne 4x\cdot 4x\] \[\huge 4x\cdot 4x=(4x)^2\ne4x^2\]

OpenStudy (anonymous):

Would the answer be 44x^5?

OpenStudy (anonymous):

Oh okay!

OpenStudy (anonymous):

@PeterPan would the answer be 44x^5?

OpenStudy (anonymous):

No... I told you to distribute 4x²(2x + 9) what do you get?

OpenStudy (anonymous):

well @PeterPan